OpenStudy (anonymous):

Given X/12=5/4 , which is proportional to x+5/x-5 A 2/1 B 1/2 C 4/5 D 5/4

OpenStudy (campbell_st):

from the 1st x/12 = 5/4 solve for x x = 60/4 x=15 then the 2nd (x + 5)/(x - 5) becomes (15 + 5)/(15 -5) = 2/1

OpenStudy (anonymous):

thank you!